OpenStudy (anonymous):

simplify the rational expressions. state any excluded values. show work. 3x-6/x-2

OpenStudy (cggurumanjunath):

(3x-6)/(x-2)=3(x-2)/(x-2)=3

OpenStudy (anonymous):

wow... thanks man.. i got one more but ill do it seperate so i can reward u!